### v3

# Estimating Interpretable Heterogeneous Treatment Effect with Causal Subgroup Discovery in Survival Outcomes

Abstract
Estimating heterogeneous treatment effects (HTE) for survival outcomes has gained increasing attention in precision medicine, as it captures the variation in treatment efficacy across patients or subgroups in disease management. However, most existing methods prioritize post-hoc subgroup identification rather than simultaneously estimating HTE and identifying causal subgroups. In this paper, we propose an interpretable HTE estimation framework that integrates meta-learners with tree-based methods to estimate the conditional average treatment effect (CATE) for survival outcomes and identify predictive subgroups simultaneously. We evaluated the performance of our method through comprehensive simulation studies. We also demonstrated its application in a large randomized controlled trial (RCT) for age-related macular degeneration (AMD), a polygenic, progressive eye disease, to estimate the HTE of an antioxidant and mineral supplement on time-to-AMD progression and to identify genetically defined subgroups with enhanced treatment effects. Our method offers a direct interpretation of the estimated HTE and provides evidence to support precision healthcare.
Keywords: interpretable causal analysis, heterogeneous treatment effect, precision medicine, subgroup identification, survival data
1 Introduction
Estimating heterogeneous treatment effects (HTE) has gained significant attention in healthcare research, as it captures variations in treatment efficacy between individuals or subgroups exposed to the same treatment. In particular, there is increasing interest in estimating HTE for survival outcomes, which provides a deeper understanding of treatment heterogeneity in disease progression and helps to develop targeted therapeutic strategies tailored to patient-specific characteristics.
HTE is assessed by estimating the conditional average treatment effect (CATE), defined as the expected difference between potential outcomes in the treatment and control arms, given patient characteristics. Existing methods for CATE estimation in survival outcomes primarily focus on the accurate estimation of CATE and post-hoc subgroup identification. For instance, Bo et al. (2024, 2025) employed meta-learners with machine learning to estimate CATE for survival data, identifying beneficial subgroups by selecting thresholds on predicted CATEs and applying variable importance techniques post hoc. Hu et al. (2021, 2022) used Bayesian methods to estimate CATE on the median survival time scale, identifying subgroups by examining posterior average survival effects against covariates. Cui et al. (2023) introduced causal survival forests (CSF) to improve the interpretability of CATE estimates through tree structures, although subgroup identification remained challenging due to the complexity of the ensemble trees.
These methods utilizing machine learning techniques often lack interpretability due to their “black box” nature, limiting their interpretations in clinical practice. To address this challenge, there is a growing need for interpretable approaches to CATE estimation. For example, Bargagli-Stoffi et al. (2024) proposed a causal rule ensemble method to estimate CATE for continuous outcomes. Their two-step approach begins by generating potential subgroups using tree-based methods, where predicted CATEs using existing causal inference methods (e.g., causal forests) are used as outcomes and regressed on covariates. In the second step, a penalized regression is applied, where the predicted CATEs are regressed on the generated subgroups. Wu et al. (2023) proposed a similar two-step approach, generating potential subgroups from causal forests and then estimating CATE through D-learner using the generated subgroups as new covariates. Wan et al. (2023a) proposed an interpretable CATE estimation method for continuous outcomes, focusing on the main effects of treatment. They first generated subgroups by regressing inverse probability-weighted outcomes on covariates using gradient-boosting trees. Then, they employed an adaptive group lasso to fit observed outcomes to these subgroups. Hiraishi et al. (2024) extended it to incorporate prognostic effects.
Interpretable CATE estimation in survival outcomes remains under-explored. Only recently, Wan et al. (2023b) extended their previous work (Wan et al., 2023a) by using a similar approach based on Cox proportional hazards (Cox PH) model. In this paper, we propose an interpretable HTE estimation framework to estimate CATE in survival outcomes, which not only predicts CATE but also identifies important subgroups. Our study is motivated by age-related macular degeneration (AMD), a progressive polygenic neurodegenerative eye disease that can cause blindness in older individuals. Randomized clinical trials (RCTs), known as the Age-Related Eye Diseases Study (AREDS and AREDS2), evaluated the efficacy of the AREDS formula (an antioxidant and mineral supplement) in delaying AMD progression, but showed minimal overall effects. Bo et al. (2024) estimated the CATE of the AREDS formula on delaying time-to-AMD progression and identified key Single Nucleotide Polymorphisms (SNP) using the post-hoc algorithm based on predicted CATEs. However, this approach only considered SNPs individually, ignoring potential combined effects. In this study, we leverage the interpretability of tree nodes from tree-based algorithms to develop an interpretable framework to estimate CATE in survival outcomes. Instead of using the Cox PH model, which relies on the PH assumption as Wan et al. (2023b) did, we employ inverse probability censoring weighting (IPCW) to handle censored subjects. Our proposed framework utilizes the idea of “pseudo-individualized treatment effect” (pseudo-ITE) (Bo et al., 2025) to address the fundamental causal inference problem, where only one potential outcome is observable. This framework integrates three meta-learners with different statistical properties to construct pseudo-ITE. We evaluated their performance in terms of prediction and subgroup identification accuracy through comprehensive simulation studies, providing practical recommendations for implementation. Finally, we applied our proposed method to the AREDS trial to estimate CATE and identify SNP-based subgroups.
The paper is organized as follows. In Section 2, we introduce the notation, assumptions, and target estimand. In Section 3, we describe the proposed framework. In Section 4, we examine the proposed methods through comprehensive simulations. A real data application is presented to illustrate the proposed method in Section 5. Finally, we conclude and discuss limitations and future work in Section 6.
2 Notation, assumptions and causal estimand
HTE is defined through the CATE function. We follow the Neyman–Rubin counterfactual framework (Rubin, 1974; Splawa-Neyman et al., 1990) to define the CATE for survival outcomes. Consider a study of sample size $n$, which compares two treatment arms, $A_{i}\in\{0,1\}$ where $A_{i}=1$ denotes taking the treatment and $A_{i}=0$ denotes taking the control for subject $i$, $i\in\mathcal{I}=\{1,2,\cdots,n\}$. $\boldsymbol{X}_{i}\in\mathbb{R}^{p}$ is a $p-$dimensional vector of covariates. Let $T$ denote the survival time and $C$ denote the censoring time. $T_{i}(1)$ and $T_{i}(0)$ denote potential survival times; $C_{i}(1)$ and $C_{i}(0)$ denote potential censoring times; $\delta_{i}(1)$ and $\delta_{i}(0)$ denote potential event indicators. We define $U=\min(T,C)$ as the observed time with the event indicator $\delta$ defined as $\delta=I(T<C)$. The observed data are denoted as $D=\{(U_{i},\delta_{i}),A_{i},\boldsymbol{X}_{i};i=1,\dots,n\}$. In this paper, we define CATE as the expected difference between $I(T(1)>t^{*})$ and $I(T(0)>t^{*})$ given covariates:
$\tau(\boldsymbol{x};t^{*})=\mathbb{E}[I(T(1)>t^{*})-I(T(0)>t^{*})|\boldsymbol{X}=\boldsymbol{x}],$ (1)
where $t^{*}$ denotes a pre-specified time of interest. CATE can also be defined for other survival quantities. For example, on the restricted mean survival time scale, CATE is defined as $\tau(\boldsymbol{x};t^{*})=\mathbb{E}[\min(T(1),t^{*})-\min(T(0),t^{*})|\boldsymbol{X}=\boldsymbol{x}]$, where $t^{*}$ is the pre-specified restricted time.
In this paper, we focus on CATE under the survival probability scale, defined in Equation (1). We assume the following causal analysis assumptions:
  • Assumption 1 (Consistency): $T_{i}=A_{i}T_{i}(1)+(1-A_{i})T_{i}(0)$, $C_{i}=A_{i}C_{i}(1)+(1-A_{i})C_{i}(0)$, $\delta_{i}=A_{i}\delta_{i}(1)+(1-A_{i})\delta_{i}(0)$
  • Assumption 2 (Unconfoundedness): $A_{i}\!\perp\!\!\!\perp(T_{i}(0),T_{i}(1))|\boldsymbol{X}_{i}$,
    $A_{i}\!\perp\!\!\!\perp(C_{i}(0),C_{i}(1))|\boldsymbol{X}_{i}$
  • Assumption 3 (Population Overlap): $e(\boldsymbol{x})=P(A_{i}=1|\boldsymbol{X}_{i}=\boldsymbol{x})\in(0,1)$
where $e(\boldsymbol{x})$ is a propensity score for treatment assignment. Given Assumptions 1 to 3, the causal estimand in Equation (1) can be identified as
$\tau(\boldsymbol{x};t^{*})$ $=\mathbb{E}[I(T>t^{*})|\boldsymbol{X}=\boldsymbol{x},A=1]-\mathbb{E}[I(T>t^{*})|\boldsymbol{X}=\boldsymbol{x},A=0]$ (2)
$=S_{1}(t^{*}|\boldsymbol{x})-S_{0}(t^{*}|\boldsymbol{x})$
where $S_{0}(t^{*}|\boldsymbol{x})$ and $S_{1}(t^{*}|\boldsymbol{x})$ are conditional survival probabilities under control and treatment, given covariates. We also assume the noninformative censoring given covariates and treatment: $T_{i}(a)\!\perp\!\!\!\perp C_{i}(a)|\{\boldsymbol{X}_{i},A_{i}\}$.
3 Methods
In this section, we introduce the proposed framework for estimating interpretable HTE in survival outcomes.
3.1 Introduction to interpretable machine learning method – RuleFit
In our proposed interpretable HTE estimation framework, we employ an interpretable machine learning algorithm called “RuleFit” (Friedman and Popescu, 2008). Below, we provide an overview of the core idea behind this algorithm.
In predictive modeling, there is often a trade-off between predictive accuracy and interpretability, particularly when using “black-box” machine learning algorithms. Consider a task where the objective is to predict a continuous outcome, such as blood pressure, based on patient covariates like age, weight, and other factors. While black-box models may yield high accuracy, they often lack transparency, limiting their clinical utility. Our goal is to develop a model that enhances interpretability without sacrificing performance. By providing clear insights into how predictions are made under specific clinical scenarios, such a model can support evidence-based decision-making and promote adoption in practice.
To enhance interpretability, the RuleFit approach constructs a prediction model of the form $y=\beta_{0}+\sum_{k=1}^{K}\beta_{k}r_{k}(\boldsymbol{x})$, where $y$ is a continuous outcome variable, $r_{k}(\boldsymbol{x})\in\{0,1\}$, $k=1,2,\cdots,K$, are binary “rule" terms, and $\beta_{k}$ are their corresponding coefficients. These rule terms represent simple logical conditions—such as whether a patient meets certain clinical criteria—and are derived from decision trees. Figure 1A gives an example in which the rule terms are generated from a classification and regression tree (CART) by regressing the outcome $y$ on the covariates. Each internal and terminal node corresponds to a binary condition that can be interpreted as a rule. From the tree of this example, we extract four rule terms. Rule 1: $r_{1}(\boldsymbol{x})=I(\text{bmi = high})$, Rule 2: $r_{2}(\boldsymbol{x})=I(\text{bmi = low})$, Rule 3: $r_{3}(\boldsymbol{x})=I(\text{bmi = high})I(\text{age}>45)$, and Rule 4: $r_{4}(\boldsymbol{x})=I(\text{bmi = high})I(\text{age}\leq 45)$. To reduce redundancy, complementary rules are removed. In this case, Rule 2 is excluded because it is complementary to Rule 1. The remaining rules (Rules 1, 3, and 4) define interpretable subgroups, allowing each subject to be evaluated for membership in these subgroups. By generating an ensemble of such trees, the covariate space $X\in\mathbb{R}^{p}$ is partitioned into binary indicators $r_{1}(\boldsymbol{x}),r_{2}(\boldsymbol{x}),\cdots,r_{K}(\boldsymbol{x})$. These indicators are then treated as new covariates in constructing an interpretable predictive model.
Figure 1: Figure A illustrates an example of candidate subgroups generated from a single decision tree. Figure B presents an example of rule extraction from an ensemble of trees. Figure C shows the flowchart of the proposed interpretable meta-learner framework for estimating CATE in survival outcomes.
In this paper, we refer to each rule as a candidate subgroup and use these candidate subgroups to construct a prediction model for estimating CATE with survival outcomes. In Section 3.2, we present the general pseudo-ITE framework (Bo et al., 2025) for estimating CATE in the context of survival data. Then, in Section 3.3, we introduce our proposed interpretable HTE estimation framework, which applies the RuleFit algorithm within the pseudo-ITE approach.
3.2 Estimating CATE through a pseudo-ITE-based meta-learner framework
As shown in Equation (1), if the ITE, defined as $I(T_{i}(1)>t^{*})-I(T_{i}(0)>t^{*})$, were observable for each subject $i$ at a pre-specified time point $t^{*}$, we could directly regress ITE on covariates to estimate CATE. However, the ITE is unobservable for two main reasons. First, even in the absence of censoring, we can only observe one of the two potential outcomes $I(T(1)>t^{*})$ or $I(T(0)>t^{*})$, since each patient receives only one treatment. Second, in the presence of censoring, $I(T>t^{*})$ cannot be determined for individuals censored before time $t^{*}$. To address these challenges, Bo et al. (2025) proposed to construct a pseudo-ITE (pITE) $Y^{*}$ based on observed data and estimate $\tau(\boldsymbol{x};t^{*})$ by minimizing a squared error-type objective function:
$\operatorname*{argmin}_{\tau}\frac{1}{n}\sum_{i=1}^{n}w_{i}^{M}\left(Y_{i}^{*}-\tau(\boldsymbol{X}_{i};t^{*})\right)^{2},$ (3)
where $Y_{i}^{*}$ is a constructed pITE that replaces the unobservable ITE; $w_{i}^{M}$ is a method-specific weight, which will be introduced in the following paragraphs.
To accommodate right-censored data, IPCW, denoted as $w^{C}=\\ \frac{1}{P(C>\min(T,t^{*})|\boldsymbol{X},A)}$, was used to predict the inverse of the probability for not being censored before $\min(T,t^{*})$ for those with $I(T>t^{*})$ being known. Thus, the final objective function to solve for CATE for right-censored data becomes:
$\operatorname*{argmin}_{\tau}\frac{1}{n^{o}}\sum_{i=1}^{n^{o}}w_{i}^{C}w_{i}^{M}\left(Y_{i}^{*}-\tau(\boldsymbol{X}_{i};t^{*})\right)^{2},$ (4)
where $n^{o}$ denotes the sample size of complete data whose $I(T>t^{*})$ is known.
There are multiple ways to construct $Y^{*}$, and in this section, we illustrate three doubly-robust approaches to construct pITE. These three methods—DR-learner, DEA-learner, and R-learner—were originally introduced in Bo et al. (2025) and are referred to as “meta-learners.” Each constructs pITE with distinct statistical properties.
DR-learner. Kennedy (2020) proposed a Doubly-Robust learner (DR-learner) derived from the efficient influence curve to estimate the average treatment effect. When formulated on the survival probability scale, the pITE can be written as follows:
$Y^{*}_{DR}=\frac{A-\hat{e}(\boldsymbol{X})}{\hat{e}(\boldsymbol{X})\left(1-\hat{e}(\boldsymbol{X})\right)}\left(I(T>t^{*})-\hat{S}_{A}(t^{*}|\boldsymbol{X})\right)+\hat{S}_{1}(t^{*}|\boldsymbol{X})-\hat{S}_{0}(t^{*}|\boldsymbol{X}).$
It satisfies $\tau(\boldsymbol{x};t^{*})=E[Y^{*}_{DR}|\boldsymbol{X}=\boldsymbol{x},t^{*}]$ and does not require a method-specific weight $w^{M}$ (or equivalently $w^{M}=1$ for all subjects).
DEA-learner. Chen et al. (2017) proposed the DEA-learner as an extension of the D-learner (Tian et al., 2014), also known as the modified covariate approach. The original objective function of D-learner solves CATE in a linear parametric form by directly minimizing an objective function, which (on the survival probability scale) can be formulated as follows:
$\operatorname*{argmin}_{\gamma}\left[\frac{1}{n^{o}}\sum_{i=1}^{n^{o}}\frac{(2A_{i}-1)\left(A_{i}-e\left(\boldsymbol{X}_{i}\right)\right)}{e\left(\boldsymbol{X}_{i}\right)\left(1-e\left(\boldsymbol{X}_{i}\right)\right)}\left(I(T_{i}>t^{*})-\frac{2A_{i}-1}{2}\boldsymbol{X}_{i}\gamma\right)^{2}\right],$
where $\tau\left(\boldsymbol{X}_{i};t^{*}\right)=\boldsymbol{X}_{i}\gamma$. To relax the restriction in linear parametric form, one can replace $\boldsymbol{X}_{i}\gamma$ with a general $\tau\left(\boldsymbol{X}_{i};t^{*}\right)$ in the objective function. Since $(2A-1)^{2}=1$ always holds for both $A=0$ and $A=1$, we can reorganize the objective function as follows:
$\operatorname*{argmin}_{\tau}\frac{1}{n^{o}}\sum_{i=1}^{n^{o}}(2A_{i}-1)\frac{A_{i}-e\left(\boldsymbol{X}_{i}\right)}{4e\left(\boldsymbol{X}_{i}\right)\left(1-e\left(\boldsymbol{X}_{i}\right)\right)}\left\{2(2A_{i}-1)I(T_{i}>t^{*})-\tau\left(\boldsymbol{X}_{i};t^{*}\right)\right\}^{2},$
where $Y_{D}^{*}=2(2A-1)I\left(T>t^{*}\right)$ is the pITE. A regression weight, $w^{M}=(2A-1)\frac{A-e\left(\boldsymbol{X}\right)}{4e\left(\boldsymbol{X}\right)\left(1-e\left(\boldsymbol{X}\right)\right)}$, needs to be applied to regress the pTIE $Y^{*}_{D}$ on covariates.
Chen et al. (2017) modified the D-learner into an efficiency augmentation version (called the “DEA-learner”) by replacing the observed outcome $I(T>t^{*})$ with its residual $I(T>t^{*})-S(t^{*}|\boldsymbol{X})$. Thus, $Y_{DEA}^{*}=\\ 2\left(2A-1\right)\left(I\left(T>t^{*}\right)-\hat{S}(t^{*}|\boldsymbol{X})\right)$ is the new pITE with the same weight $w^{M}=(2A-1)\frac{A-e\left(\boldsymbol{X}\right)}{4e\left(\boldsymbol{X}\right)\left(1-e\left(\boldsymbol{X}\right)\right)}$.
R-learner. Nie and Wager (2020) proposed the R-learner using Robinson’s decomposition: $I(T_{i}>t^{*})-E\left[I(T_{i}>t^{*})|\boldsymbol{X}_{i}=\boldsymbol{x},t^{*}\right]=\left(A_{i}-e\left(\boldsymbol{X}_{i}\right)\right)\tau\left(\boldsymbol{X}_{i};t^{*}\right)+\varepsilon_{i}$, where $E\left[\varepsilon_{i}|\boldsymbol{X}_{i},A_{i}\right]=0$. Then, CATE can be solved by minimizing the objective function:
$\operatorname*{argmin}_{\tau}\left\{\frac{1}{n^{o}}\sum_{i=1}^{n^{o}}\left[\left(I(T_{i}>t^{*})-S\left(t^{*}|\boldsymbol{X}_{i}\right)\right)-\left(A_{i}-e\left(\boldsymbol{X}_{i}\right)\right)\tau\left(\boldsymbol{X}_{i};t^{*}\right)\right]^{2}\right\},$
where $S\left(t^{*}|\boldsymbol{X}\right)$ is the survival function ignoring the treatment groups. We can rewrite the objective function as follows:
$\operatorname*{argmin}_{\tau}\frac{1}{n^{o}}\sum_{i=1}^{n^{o}}\left(A_{i}-e\left(\boldsymbol{X}_{i}\right)\right)^{2}\left\{\frac{I(T_{i}>t^{*})-S\left(t^{*}|\boldsymbol{X}_{i}\right)}{A_{i}-e\left(\boldsymbol{X}_{i}\right)}-\tau\left(\boldsymbol{X}_{i};t^{*}\right)\right\}^{2}.$
Thus, $Y_{RL}^{*}=\frac{I\left(T>t^{*}\right)-{S}(t^{*}|\boldsymbol{X})}{A-{e}(\boldsymbol{X})}$ is the pITE under this approach and $w^{M}=\left(A-{e}(\boldsymbol{X})\right)^{2}$ is the method-specific weight.
The propensity score $e(\boldsymbol{X})$ and the mean outcome regression models (i.e., $S_{A}(t^{*}|\boldsymbol{X})$ and $S(t^{*}|\boldsymbol{X})$) in the construction of $Y^{*}$ and weights $w^{M}$, $w^{C}$ are all treated as nuisance parameters and need to be estimated beforehand. Finally, the CATE $\tau(\boldsymbol{x};t^{*})$ can be estimated by minimizing the objective function (4) with $Y^{*}$, $w^{M}$ and $w^{C}$ plugged in.
3.3 Interpretable HTE estimation framework
Suppose we have $K$ candidate subgroups, $r_{1}(\boldsymbol{x}),r_{2}(\boldsymbol{x}),\cdots,r_{K}(\boldsymbol{x})$, that may potentially contribute to CATE prediction. We can re-parameterize $\tau(\boldsymbol{x};t^{*})$ in the objective function (4) as $\tau(\boldsymbol{x};t^{*})=\beta_{0}+\sum_{k=1}^{K}\beta_{k}r_{k}(\boldsymbol{x})$ for $k=1,2,\cdots,K$. Thus, our proposed interpretable framework estimates the CATE in survival outcomes by minimizing the following objective function with a penalty term to select important subgroups:
$\operatorname*{argmin}_{\boldsymbol{\beta}}\frac{1}{n^{o}}\sum_{i=1}^{n^{o}}w_{i}^{C}w_{i}^{M}\{Y_{i}^{*}-[\beta_{0}+\sum_{k=1}^{K}\beta_{k}r_{k}(\boldsymbol{x}_{i})]\}^{2}+\lambda\mathcal{R}(\boldsymbol{\beta}),$ (5)
where $r_{k}(\boldsymbol{x}_{i})\in\{0,1\}$ denotes whether a subject $i$ with covariate $\boldsymbol{x}_{i}$ falls into the $k$-th candidate subgroup or not; $\beta_{k}$ denotes the coefficient of the corresponding candidate subgroup $r_{k}(\boldsymbol{x}_{i})$; $\mathcal{R}(\boldsymbol{\beta})$ is a penalty term for subgroup selection with a positive $\lambda$ as the tuning parameter.
Algorithm 1 An interpretable HTE estimation framework for survival outcomes
Input: Covariate $\boldsymbol{X}$, treatment $A$, pre-specified time $t^{*}$, observed outcome $I(T>t^{*})$.
Step 1: Construct $Y^{*}$, $w^{M}$ and $w^{C}$
Estimate the following nuisance parameters on the training data.
(1) Estimate propensity score $e(\boldsymbol{X})$ and outcome regression models $S_{a}(t^{*}|\boldsymbol{X})$ or $S(t^{*}|\boldsymbol{X})$.
(2) Construct pITE $Y^{*}$ for individuals whose observed outcome $I(T>t^{*})$ can be determined, using the selected meta-learner.
(3) Construct the meta-learner method-specific weight $w^{M}$.
(4) Estimate IPCW $w^{C}=\frac{1}{P(C>c|\boldsymbol{X}=\boldsymbol{x},A=a)}$ using all individuals and predict $w^{C}$ for each individual whose $I(T>t^{*})$ can be determined.
Step 2. Candidate subgroup generation
(1) Minimize the objective function:
$\operatorname*{argmin}_{\tau}\frac{1}{n^{o}}\sum_{i=1}^{n^{o}}w_{i}^{C}w_{i}^{M}\left(Y_{i}^{*}-\tau(\boldsymbol{X}_{i};t^{*})\right)^{2}$
through an ensemble tree-based algorithm
(2) Extract all terminal and internal nodes from the generated trees.
(3) Remove complementary and repeated tree nodes to obtain candidate subgroups $r_{1}(\boldsymbol{x}),r_{2}(\boldsymbol{x}),\cdots,r_{K}(\boldsymbol{x})$.
Step 3. Subgroup selection
(1) Transform the original covariates $\boldsymbol{X}\in\mathbb{R}^{p}$ into binary indicators $r_{k}(\boldsymbol{x})\in\{0,1\}$, with $r_{k}(\boldsymbol{x})$ from Step 2.
(2) Minimize the following objective function through a penalized regression:
$\operatorname*{argmin}_{\boldsymbol{\beta}}\frac{1}{n^{o}}\sum_{i=1}^{n^{o}}w_{i}^{C}w_{i}^{M}\{Y_{i}^{*}-[\beta_{0}+\sum_{k=1}^{K}\beta_{k}r_{k}(\boldsymbol{x}_{i})]\}^{2}+\lambda\mathcal{R}(\boldsymbol{\beta}),$
where $\lambda$ is a tuning parameter and $\mathcal{R}(\boldsymbol{\beta})$ is a penalty term.
Output: $L$ numbers of selected subgroups ($L<K$), denoted as $r_{1}(\boldsymbol{x}),r_{2}(\boldsymbol{x}),\cdots,r_{L}(\boldsymbol{x})$, with estimated coefficients $\hat{\beta}_{l}$, $l=1,2,\cdots,L$.
For a new data point with covariate $\boldsymbol{x}$, CATE is predicted as: $\hat{\tau}(\boldsymbol{x};t^{*})=\hat{\beta}_{0}+\sum_{l=1}^{L}\hat{\beta}_{l}r_{l}(\boldsymbol{x}).$
Algorithm 1 describes the proposed interpretable HTE estimation framework for survival outcomes. Flowchart 1 visually illustrates each step of the algorithm. In Step 1, we construct $Y^{*}$, $w^{M}$, and $w^{C}$, which involves estimating the following nuisance parameters: the propensity score $e(\boldsymbol{X})$, the conditional survival function in each treatment arm $S_{a}(t^{*}|\boldsymbol{X})$) or the conditional survival function ignoring the treatment arm $S(t^{*}|\boldsymbol{X})$). Parametric or semi-parametric methods can be applied (e.g., random forests (RF) for estimating $e(\boldsymbol{X})$, random survival forests (RSF) for estimating $S_{a}(t^{*}|\boldsymbol{X})$ or $S(t^{*}|\boldsymbol{X})$). The estimated nuisance parameters are then used to construct $Y^{*}$ and $w^{M}$ within the meta-learner selected by the user. Finally, the censoring probability can be estimated using the Kaplan-Meier estimator (KM) if censoring is completely independent of survival time (i.e., $T\perp C$) or using RSF if censoring is covariate-dependent (i.e., $T\not\perp C$ but $T\perp C|\{\boldsymbol{X},A$}).
In Step 2, we plug in $Y^{*}$, $w^{C}$, and $w^{M}$ into the objective function (4) to solve for CATE using a tree-based method. This will transform the original covariate space $\boldsymbol{X}\in\mathbb{R}^{p}$ into candidate subgroups $\boldsymbol{r}(\boldsymbol{x})$. Various tree-based methods can be considered. For example, Fokkema (2020) gives examples using CART and conditional inference tree (CTree), each offering distinct advantages. CART is a well-understood and widely implemented tree-based method in many R packages; on the other hand, CTree addresses the variable selection bias by implementing unbiased splitting criteria that evaluate whether each covariate carries information about the outcome when determining the splitting point at internal nodes. Ensemble algorithms (e.g., gradient boosting, RF, etc.) can be used to extract tree nodes that collectively cover the covariate space as fully as possible. In the resulting trees, each internal or terminal node represents a potential candidate subgroup. Because complementary subgroups are perfectly correlated (i.e., $r_{k}(\boldsymbol{x})+r_{k^{\prime}}(\boldsymbol{x})=1$ for $k\neq k^{\prime}$), one subgroup from each complementary pair is removed. Additionally, duplicate subgroups generated across multiple trees are also eliminated. Figure 1B shows an example of candidate subgroups generated from trees, represented as the colored nodes.
In Step 3, we transform the original covariates into binary indicators $r_{k}(\boldsymbol{x})\in\{0,1\}$, where each $r_{k}(\boldsymbol{x})$ corresponds to a candidate subgroup generated in Step 2. These indicators $r_{1}(\boldsymbol{x}),r_{2}(\boldsymbol{x}),\cdots,r_{K}(\boldsymbol{x})$ are then used as new covariates into the objective function (5), where we solve for the coefficients $\beta$ by applying a regularization penalty term (e.g., Lasso, Elastic Net) with cross-validation (CV) to select the tuning parameter $\lambda$. For example, Friedman and Popescu (2008) implemented the Lasso penalty, which yields a sparse set of selected subgroups and improves interpretability. In this paper, we also adopt the Lasso penalty to enhance the interpretability of HTE prediction. The resulting objective function in Step 3 becomes:
$\operatorname*{argmin}_{\boldsymbol{\beta}}\frac{1}{n^{o}}\sum_{i=1}^{n^{o}}w_{i}^{C}w_{i}^{M}\{Y_{i}^{*}-[\beta_{0}+\sum_{k=1}^{K}\beta_{k}r_{k}(\boldsymbol{x}_{i})]\}^{2}+\lambda\sum_{k=1}^{K}|\beta_{k}|.$ (6)
The predicted CATE is $\hat{\tau}(\boldsymbol{x};t^{*})=\hat{\beta}_{0}+\sum_{l=1}^{L}\hat{\beta}_{l}r_{l}(\boldsymbol{x})$, where $r_{l}(\boldsymbol{x})$’s are final selected subgroups ($L<K$), determined based on the tuning parameter $\lambda$ value, which is choosen to make the best predictions through the CV.
4 Simulations
In this section, we present comprehensive simulation studies conducted under various study designs, including both low-dimensional independent settings and high-dimensional correlated signal settings.
4.1 Simulation designs
We first considered an independent signal setting with ten independent covariates. $X_{1},X_{2},\cdots,X_{5}$ were binary covariates created by $X_{j}=I(\tilde{X}_{j}>0)$, where $\tilde{X}_{j}$ was independently generated from a normal distribution $N(0,1)$, and $X_{6},X_{7},\cdots,X_{10}$ were continuous covariates independently generated from $N(0,1)$. Treatment assignment was determined by the propensity score: $\mathrm{logit}(e(\boldsymbol{X}))=0.4-0.3X_{1}-0.2X_{6}-0.3X_{2}-0.35X_{7}-0.2X_{3}-0.25X_{8}$.
Potential survival times were simulated from a Weibull regression model: $T(a|\boldsymbol{X})=\lambda_{a}\{\frac{-\mathrm{log}(U)}{\mathrm{exp}\{f_{a}(\boldsymbol{X})\}}\}^{\frac{1}{\eta}}$, where $U\sim\mathrm{Unif}[0,1]$ and $a=0$ or $1$. $f_{a}(\boldsymbol{X})=b(\boldsymbol{X})+a*h(\boldsymbol{X})$ introduced the source of HTE. $b(\boldsymbol{X})$ was a function of covariates shared between two treatment arms; $h(\boldsymbol{X})$ is another function of covariates only for the treatment arm. We considered simple to complex scenarios of $b(\boldsymbol{X})$ and $h(\boldsymbol{X})$ listed below. In Scenario 1, the covariates are in the linear form and do not share between $b(\boldsymbol{X})$ and $h(\boldsymbol{X})$. In Scenario 2, we added one shared covariate in both $b(\boldsymbol{X})$ and $h(\boldsymbol{X})$. In Scenario 3, we considered the nonlinear form and/or interaction of the covariates in both $b(\boldsymbol{X})$ and $h(\boldsymbol{X})$. Under this setting, around half of the covariates are independent noise variables that do not contribute to either $b(\boldsymbol{X})$ or $h(\boldsymbol{X})$. Note that the subgroups are defined based on covariates in $b(\boldsymbol{X})$ and $h(\boldsymbol{X})$, where covariates in $h(\boldsymbol{X})$ directly contribute to CATE predictions, while those in $b(\boldsymbol{X})$ contribute indirectly. The shape parameter was set as $\eta=2$; scale parameter was set as $\lambda_{0}=16,20,20$ and $\lambda_{1}=26,22,22$ for scenarios 1 to 3 respectively. The censoring time was simulated independently from an exponential distribution to yield a 30% censoring rate for each scenario.
[…]
$h(\boldsymbol{X})$ $=$ $2.8X_{1}+1.4X_{2}.$
[…]
$h(\boldsymbol{X})$ $=$ $2.5X_{1}-1.8X_{2}-2X_{3}.$
[…]
$h(\boldsymbol{X})$ $=$ $2.5X_{1}-1.8X_{2}-2X_{3}-1.4X_{1}X_{3}.$
Additionally, we considered high-dimensional settings with correlated signals by generating covariates from a multivariate normal distribution $\tilde{\boldsymbol{X}}\sim MVN(\boldsymbol{0},0.5\Sigma)$, where $\Sigma=\{\sigma_{jj^{\prime}}=e^{-|j-j^{\prime}|},1\leq j,j^{\prime}\leq p\}$ for $p=100$. $X_{1},X_{2},X_{3},X_{4},X_{5},X_{11},X_{12},\cdots,X_{55}$ are binary covariates created by $X_{j}=I(\tilde{X}_{j}>0)$ for $j=1,2,3,4,5,11,12,\cdots,55$ and the rest are continuous with $X_{j}=\tilde{X}_{j}$. We generated 100 training datasets of sample size $n=1,000$ and one test dataset of sample size $N=10,000$.
In Step 1, different machine learning methods can be applied to estimate conditional survival functions or the propensity score. Bo et al. (2024) studied the performance of different machine learning methods, including RSF, Bayesian accelerated failure time model (BAFT) (Henderson et al., 2018) and deep survival neural network (DNNSurv) (Sun et al., 2020), to estimate conditional survival functions $S_{a}(t|\boldsymbol{x})$. With a moderate to large sample size, RSF achieves satisfactory predictive performance while maintaining fast computational speed. In this paper, we applied RSF (Ishwaran and Kogalur, 2007) to estimate $S_{a}(t^{*}|\boldsymbol{X})$ and $S(t^{*}|\boldsymbol{X})$ and RF (Liaw and Wiener, 2002) to estimate $e(\boldsymbol{X})$. Out-of-bag prediction was used to predict the propensity score and conditional survival functions for each patient. In this simulation, we used RF to estimate $e(\boldsymbol{X})$ and used the KM method to estimate $P(C>\min(T,t^{*}))$ under this independent censoring setting. If censoring is potentially dependent on covariates, we recommend using model-based or machine learning approaches (e.g., RSF) to estimate the censoring probability. In Section 4.3.3, we also evaluate the performance of our method under high censoring rates (e.g., 50% to 70%) and covariate-dependent censoring scenarios.
In Step 2, we implemented CTree with a gradient boosting algorithm to generate an ensemble of shallow trees through the R package pre as CTree implements unbiased splitting criteria (Friedman and Popescu, 2008) that enhances the ability of trees to split on covariates that truly contribute to CATE. Specifically, it tests the hypothesis $H^{j}_{0}:P(Y^{*}|\boldsymbol{X}_{j})=P(Y^{*})$ with the global null hypothesis $H_{0}:\bigcap_{j=1}^{p}H_{0}^{j}$. The partial hypothesis is tested through permutation tests. Multiple testing adjustments can be applied (e.g., Bonferroni corrections). The tree stops splitting if the global null hypothesis is not rejected at a pre-specified $\alpha$ level. Note that CTree is already unbiased splitting without multiple testing adjustments. More details can be found in their original paper (Hothorn et al., 2006). The maximum tree depth was set as 5, with a maximum of 500 trees to be generated. In the independent signal setting, Bonferroni correction in CTree was implemented with $\alpha=0.1$. No multiple testing adjustment was conducted for the high-dimensional correlated signal setting.
In Step 3, we used the Lasso penalty and performed a 10-fold CV to select the tuning parameter $\lambda$ that yields the best predictions in terms of the smallest mean squared errors and output finally selected subgroups $r_{1}(\boldsymbol{x}),r_{2}(\boldsymbol{x}),\cdots,r_{L}(\boldsymbol{x})$.
Finally, we calculated subgroup importance score and variable importance score, denoted as: $R_{l}=|\hat{\beta}_{l}|\cdot\sqrt{s_{l}(1-s_{l})}$ and $V_{j}=\sum_{X_{j}\in r_{l}}\frac{R_{l}}{c_{l}}$, where $s_{l}=\frac{1}{n^{o}}\sum_{i=1}^{n^{o}}r_{l}(\boldsymbol{x}_{i})$ denotes the support of subgroup $r_{l}$ and $c_{l}$ denotes the number of subgroups that contain $X_{j}$. We also evaluated the prediction performance on test data based on the metrics introduced in Section 4.2.
We compared the performance of DR-, DEA-, and R-learners with the scenarios when pITE $Y^{*}$ was generated directly using the predicted $\hat{\tau}(\boldsymbol{x};t^{*})$ from CSF (Cui et al., 2023) or BAFT. For CSF, we set the number of trees to 500, mtry $=p$, and the minimum node size to 15. For BAFT, we generated 500 trees with 100 MCMC iterations and obtained 1,000 posterior draws. In this paper, we estimated CATE at the median survival time (i.e., $t^{*}$ = median survival time). One may select different times of interest depending on the specific research question.
4.2 Evaluation metrics
Prediction performance. We evaluated the prediction performance of each method in terms of Bias, binned-RMSE, and Spearman rank correlation, defined as:
Bias $=$ $\frac{1}{N}\sum\limits_{i=1}^{N}(\hat{\tau}_{i}(\boldsymbol{x};t^{*})-\tau_{i}(\boldsymbol{x};t^{*})),\quad$
binned-RMSE $=$ $\frac{1}{Q}\sum\limits_{q=1}^{Q}\sqrt{\frac{1}{N_{q}}\sum\limits_{i=1}^{N_{q}}(\hat{\tau}_{i}(\boldsymbol{x};t^{*})-\tau_{i}(\boldsymbol{x};t^{*}))^{2}},\quad$
Spearman Corr $=$ $1-\frac{6\sum d_{i}^{2}}{N(N^{2}-1)},$
where $\tau_{i}(\boldsymbol{x};t^{*})$ and $\hat{\tau}_{i}(\boldsymbol{x};t^{*})$ denote the true and predicted CATE value for subject $i$, respectively, and $d_{i}$ denotes the difference between the ranks of $\tau_{i}(\boldsymbol{x};t^{*})$ and $\hat{\tau}_{i}(\boldsymbol{x};t^{*})$ in the testing data. To evaluate the treatment heterogeneity, we evaluated CATE prediction within each subgroup using the binned-RMSE (Hu et al., 2021). We divided the predictions into $Q$ bins based on the ordering of true CATEs, calculated the MSE within each bin, and then averaged across all bins. We set $Q=50$ here.
Subgroup identification accuracy. We evaluated each method’s ability to select the correct covariates that contribute to CATE prediction in either $b(\boldsymbol{X})$ or $h(\boldsymbol{X})$. Specifically, we recorded how often each covariate was used to define at least one selected subgroup across 100 simulation runs. In addition, we examined the total number of selected subgroups across these 100 simulations.
4.3 Simulation results
4.3.1 Independent signal setting
Figure 2 shows prediction performance under the independent signal setting with Bonferroni corrections ($\alpha=0.1$) in CTree for generating candidate subgroups. In scenario 1, DEA-learner and BAFT produce minimal biases compared to other methods. In Scenarios 2 and 3, biases increase slightly for all methods except BAFT, which remains the smallest bias. Using predicted CATEs from CSF shows the largest bias, particularly in complex scenarios (2 and 3). All methods show comparable RMSEs and high correlations with true CATEs, where BAFT slightly outperforms the others. Overall, when using three meta-learners to construct $Y^{*}$ under our framework, they show comparable prediction accuracy to the case when $Y^{*}$ is generated as predicted CATE from BAFT.
Figure 2: Prediction accuracy under the independent signal setting with 30% independent censoring. The prediction performance under the following meta-learners was examined: DR-learner (DR), R-learner(R), and DEA-learner (DEA). “CSF” represents causal survival forests. “BAFT” represents the Bayesian accelerated failure time model. Bonferroni correction with $\alpha=0.1$ was performed in generating candidate subgroups using gradient boosting CTree. The upper panel shows biases, the middle panel shows binned-RMSEs, and the lower panel shows Spearman rank correlations.
Figure 3 displays the frequency with which each covariate is included in at least one selected subgroup across 100 simulations. Under the DEA-learner, covariates in $h(\boldsymbol{X})$ (those directly contributing to CATE) are frequently selected to form subgroups (100% for $X_{1}$, $>60\%$ for $X_{2}$), while covariates in $b(\boldsymbol{X})$ appear in about 40% of simulations. Noise covariates (i.e., those not belonging to $h(\boldsymbol{X})$ or $b(\boldsymbol{X})$) are rarely selected with frequencies below 10% across all three scenarios. The DR-learner tends to select noise covariates more frequently in all scenarios. Although BAFT achieves strong predictive performance, it includes all noise covariates in every simulation. The R-learner occasionally fails to generate more than one candidate subgroup, resulting in no final subgroups being identified in some runs. Thus, the frequencies of covariates used in generating subgroups under the R-learner are lower than in other methods.
Figure 3: Variable detection under the independent signal setting with 30% independent censoring. The histogram shows the number of times each covariate is used to formulate any selected subgroups among 100 simulations. Purple represents covariates in $h(\boldsymbol{X})$; orange represents covariates in $b(\boldsymbol{X})$; green represents covariates shared between $h(\boldsymbol{X})$ and $b(\boldsymbol{X})$; grey represents noise covariates appearing in neither $b(\boldsymbol{X})$ nor $h(\boldsymbol{X})$. The prediction performance under the following meta-learners was examined: DR-learner (DR), R-learner(R), and DEA-learner (DEA). “CSF” represents causal survival forests. “BAFT” represents the Bayesian accelerated failure time model. Bonferroni correction with $\alpha=0.1$ was performed to generate candidate subgroups using gradient boosting CTree.
The upper panel of Table 1 summarizes the number of selected subgroups under each method. DEA-learner retains the sparseness, selecting 3 to 10 median number of subgroups among 100 simulations. DR-learner also retains some sparseness but selects more subgroups than the DEA-learner (the median number of selected subgroups is between 11 and 24). BAFT and CSF select more than 100 subgroups that contain a lot of noise covariates. In addition, the effects of these selected subgroups under BAFT and CSF are minimal (i.e., $\hat{\beta}_{k}<0.01$ for most coefficients).
Table 1: The number of selected subgroups under each method in the independent signal setting (Bonferroni correction $\alpha=0.1$ in gradient boosting CTree), and the high-dimensional correlated signal setting ($\alpha=0.01$ without Bonferroni correction in gradient boosting CTree) among 100 simulations, median (min, max).
DR R DEA BAFT CSF
Independent Signals
Scenario 1 11 (0, 35) 0 (0, 6) 3 (0, 10) 153 (81, 203) 124 (72, 182)
Scenario 2 24 (4, 65) 3 (0, 8) 10 (3, 19) 227 (166, 295) 164 (106, 281)
Scenario 3 21 (0, 63) 2 (0, 6) 7 (1, 14) 179 (136, 231) 138 (96, 239)
High-dimensional Correlated Signals
Scenario 1 585 (403, 694) 0 (0, 4) 24 (3, 43) 460 (349, 530) 383 (303, 616)
Scenario 2 623 (458, 735) 2 (0, 7) 26 (5, 53) 535 (418, 612) 528 (368, 721)
Scenario 3 609 (435, 716) 2 (0, 5) 28 (3, 60) 480 (392, 588) 449 (312, 648)
4.3.2 High-dimensional correlated signal setting
Figure 4 presents the prediction accuracy under the correlated signal setting with 30% independent censoring. Unlike the low-dimensional independent signal setting, $\alpha$ was set as 0.01 without applying Bonferroni corrections when generating candidate subgroups using CTree. In high-dimensional and correlated settings, applying Bonferroni corrections—even with a larger $\alpha$— may result in no candidate subgroups being selected. Consistent with the independent signal setting, both DR- and DEA-learners exhibit minimal bias in Scenario 1, with slightly increased bias in Scenarios 2 and 3. As in the low-dimensional setting, BAFT yields the smallest biases across all three scenarios. In contrast, CSF shows larger biases than the other methods, particularly in Scenarios 2 and 3. DR-learner also produces the highest RMSEs and the lowest correlations with the true CATE values across all scenarios, while the remaining methods demonstrate comparable RMSEs and correlations.
Figure 4: Prediction accuracy under the high-dimensional correlated signal setting with a 30% independent censoring rate. The prediction performance under the following meta-learners was examined: DR-learner (DR), R-learner (R), and DEA-learner (DEA). “CSF” represents causal survival forests. “BAFT” represents the Bayesian accelerated failure time model. $\alpha$ was set as 0.01 without Bonferroni corrections for generating candidate subgroups using gradient boosting CTree. The upper panel shows biases, the middle panel shows binned-RMSEs, and the lower panel shows Spearman rank correlations.
Figure 5 shows the number of times each covariate is used to formulate any selected subgroup across 100 simulations. Similar to the independent signal setting, the DEA-learner selects covariates in $h(\boldsymbol{X})$ to form subgroups in over 80% of simulations. In contrast, noise covariates are infrequently used, with low selection frequencies (around 20%) across all scenarios. The DEA-learner also selects covariates in $b(\boldsymbol{X})$ with moderate frequencies. In comparison, the DR-learner, BAFT, and CSF include a substantial number of noise covariates when forming subgroups. Notably, the R-learner only generates one candidate subgroup in 80%, 40%, and 30% simulations from Scenarios 1 to 3, respectively. These runs were excluded from evaluation since the Lasso penalty requires at least two covariates as input. While the R-learner rarely selects noise variables, it also fails to consistently identify signal variables. As shown in the lower panel of Table 1, the DEA-learner maintains sparsity in the selected subgroups, whereas the DR-learner, BAFT, and CSF each yield more than 380 subgroups with small effect sizes. The R-learner fails to generate subgroups in many simulation runs, with a median number of selected subgroups ranging from 0 to 2.
Figure 5: Variable detection under the high-dimensional correlated signal setting with 30% independent censoring. The histogram shows the number of times each covariate is ever used to formulate any selected subgroups among 100 simulations. Purple represents covariates in $h(\boldsymbol{X})$; orange represents covariates in $b(\boldsymbol{X})$; green represents covariates shared between $h(\boldsymbol{X})$ and $b(\boldsymbol{X})$; grey represents noise covariates that are not in either $b(\boldsymbol{X})$ or $h(\boldsymbol{X})$. The prediction performance under the following meta-learners was examined: DR-learner (DR), R-learner (R), and DEA-learner (DEA). “CSF” represents causal survival forests. “BAFT” represents the Bayesian accelerated failure time model. $\alpha$ was set as 0.01 without Bonferroni corrections for generating candidate subgroups using gradient boosting CTree.
4.3.3 Sensitivity analysis of different censoring mechanisms
We also conducted simulations considering different censoring situations for the independent signal setting. First, we varied the censoring rate. Web Figure 1 shows the prediction accuracy with 60% censoring under the independent signal setting, while Web Figure 2 presents the variable detection results. The upper panel of Web Table 1 displays the number of selected subgroups. We observe similar prediction accuracy as seen in the low censoring rate setting. However, the DEA-learner, which typically shows low frequencies in selecting noise variables, demonstrates higher frequencies in selecting noise variables with 60% censoring (Web Figure 2). The DEA-learner selected a median of 7 to 14 subgroups after Lasso penalization, which is slightly more than in the low censoring rate setting. Other methods perform similarly to the low censoring setting.
We then considered a covariate-dependent censoring setting. Specifically, censoring times were generated from a Weibull regression model: $C=\frac{-2log(U)}{\exp(b_{0}+b_{1}X_{1}+b2X_{2})}$, where $U\sim\mathrm{Unif}[0,1]$, and values of $b$’s are chosen to yield a 30% censoring rate for Scenarios 1 to 3 respectively. We observe similar prediction accuracy (Web Figure 3), variable selection patterns (Web Figure 4), and numbers of selected subgroups (middle panel of Web Table 1) compared to the independent censoring setting with the same censoring rate.
In summary, across both low-dimensional independent and high-dimensional correlated signal settings, the DEA-learner consistently performs best in selecting sparse subgroups that include the correct covariates contributing to CATE, while also maintaining strong predictive accuracy. However, it tends to select more noise variables when the censoring rate is high. Notably, in high-dimensional, correlated settings, applying Bonferroni corrections in CTree results in approximately 5% of simulations with no subgroups selected, even though the DEA-learner still demonstrates the best overall performance (see supplemental materials for details). Therefore, in high-dimensional correlated scenarios, we recommend avoiding multiple testing corrections and instead using a smaller significance threshold (e.g., $\alpha=0.01$).
5 Real data application
5.1 Data description
In this study, we analyzed data from two clinical trials, AREDS and AREDS2. The AREDS trial served as our training dataset, and the subsequent AREDS2 trial served as the test data. We analyzed 806 patients in AREDS data who were diagnosed with AMD but free of late AMD in at least one eye at enrollment, categorized as AMD stages 2, 3, or 4. As shown in Table 2, patients were randomized based on their baseline disease severity and followed for up to 12 years, with 415 (51.49%) randomized to the AREDS formula arm (treatment) and 391 (48.51%) to the placebo arm (control). The study population had a mean age of 68.77 years (SD = 5.05) at enrollment, 466 (57.82%) female, and 393 (48.76%) non-smokers. Patients included in AREDS had mild-to-moderate AMD, with approximately half (56.70%) of the patients having moderate AMD and a mean baseline AMD severity scale of 4.09 (SD = 2.06). In addition to the baseline characteristics, we included 46 SNPs (denoted as “CE4” SNPs), which were identified to be associated with treatment efficacy by the CE4 method (Wei et al., 2021) and 67 prognostic SNPs, which were reported to be associated with AMD progression (Yan et al., 2018) (top SNPs with $p<10^{-5}$).
Table 2: Baseline characteristics of the AREDS and AREDS2 data used in the analysis.
AREDS trial AREDS2 trial
Number of subjects
All
($n=806$)
Placebo
($n=391$)
AREDS formula⋆⋆
($n=415$)
p-value
AREDS formula ⋆⋆
($n=112$)
Age 0.4905
Mean (SD) 68.77 (5.05) 68.90 (5.17) 68.66 (4.93) 71.29 (7.25)
Median (Range) 68.60 (55.30-81.00) 68.50 (55.30-81.00) 68.70 (55.50-79.50) 71.00 (53.00, 86.00)
Sex (n, %) 0.8236
Female 466 (57.82) 224 (57.29) 242 (58.31) 50 (44.64)
Male 340 (42.18) 167 (42.71) 173 (41.69) 62 (55.36)
Education (n, %) 0.3702
$\leq$ High School 269 (33.37) 124 (31.71) 145 (34.94) 27 (24.32)
College or above 537 (66.63) 267 (68.28) 270 (65.06) 84 (75.68)
Smoking (n, %) 0.6877
Never Smoked 393 (48.76) 194 (49.62) 199 (47.95) 44 (39.29)
Former/Current Smoker 413 (51.24) 197 (50.38) 216 (52.05) 68 (60.71)
Baseline AREDS AMD severity score 0.6303
Mean (SD) 4.09 (2.06) 4.13 (2.06) 4.06 (2.07) 5.43 (0.96)
$p$-value is based on two-sample $t$-test or Chi-square test for continuous or categorical variables in AREDS trial.  ⋆⋆ AREDS formula: Antioxidants and Zinc
AREDS2 was a subsequent clinical trial that enrolled patients with more advanced AMD and followed them for six years. It evaluated the efficacy of four treatment arms—modifications of the original AREDS formula, one of which was the original formula itself. For this analysis, we included 112 patients from the AREDS formula arm who had comparable AMD severity at baseline to the AREDS participants. These patients had a mean baseline severity score of 5.43 (SD = 0.96), a mean enrollment age of 71.29 years (SD = 7.25), with 50 (44.64%) identifying as female and 44 (39.29%) as non-smokers (see right column of Table 2). After excluding two patients with missing SNP data, 110 patients were included in the final analysis.
5.2 Interpretable prediction of CATE of AREDS formula on time-to-AMD progression
In this section, we applied the proposed interpretable HTE estimation framework to the AREDS clinical trial to estimate the CATE for time-to-AMD progression. The time point of interest was five years, a clinically meaningful milestone also used in Bo et al. (2024). We used the DEA-learner to construct pseudo-ITE, as it demonstrated the best performance under high-dimensional correlated signal settings in our simulation studies. RSF (Ishwaran and Kogalur, 2007) and RF (Liaw and Wiener, 2002) were performed to estimate $S_{A}(t^{*}|\boldsymbol{X})$ and $e(\boldsymbol{X})$, respectively, using the same tuning parameter setup as in the simulations. The model was trained on the AREDS dataset and evaluated on the AREDS2 dataset. To generate candidate subgroups, we implemented CTree with a gradient boosting algorithm, consistent with the approach used in the simulations. We set $\alpha=0.01$ without Bonferroni adjustment, following the same parameter settings and recommendations derived from our simulation results.
Compared to previous studies on AMD (Wei et al., 2021; Yan et al., 2018; Bo et al., 2025), our approach identified subgroups that interpret CATE predictions by capturing interaction effects between SNPs and/or baseline covariates. In contrast, prior work primarily focused on identifying individual SNPs through genome-wide association studies or post-hoc variable selection following CATE estimation using “black-box” machine learning methods. Figure 6 presents the subgroups identified by training the DEA-learner on the AREDS dataset using our proposed interpretable HTE estimation framework for survival outcomes. A total of nineteen subgroups were found to be predictive of CATE in delaying time-to-AMD progression. Fourteen SNPs and the baseline AMD severity scale formulate these nineteen subgroups, in which nine SNPs are CE4 SNPs (i.e., associated with treatment efficacy) with one from CHR 3 ($rs$9815579), one from CHR 5 ($rs$149309589), three from CHR 10 ($rs$1618927,$rs$1871453, $rs$4747238), three from CHR 14 ($rs$1056437, $rs$4903476, $rs$77000175) and one from CHR 19 ($rs$141380308); and five SNPs are prognostic SNPs (i.e., associated with disease progression regardless of the treatment), including $rs$59182762 (CHR 1), $rs$11132213 (CHR 4), $rs$79069165 (CHR 5), $rs$665731 (CHR 11), $rs$12930861 (CHR 16). Among the nineteen identified subgroups, twelve are defined by interactions between two or more covariates. Subgroups with positive estimated coefficients (highlighted in orange) represent beneficial subgroups, indicating enhanced treatment effects. In contrast, subgroups with negative estimated coefficients (highlighted in blue) are considered adverse subgroups, where treatment effects are diminished or harmful. The complement of these adverse subgroups can also be interpreted as beneficial. The right column of Figure 6 displays the size of each subgroup within the AREDS trial population. It is important to note that the selected subgroups are not mutually exclusive. The primary goal of our proposed method is to enhance the interpretability of CATE predictions by uncovering interaction effects among covariates, rather than to identify all disjoint subgroups. External validation is necessary to assess the reproducibility of the identified subgroups. In the following section, we validate these results using the AREDS2 dataset.
Figure 6: Subgroups identified by training the algorithm on AREDS at year five. The right panel shows the subgroup importance ordered by the subgroup importance score. Subgroup size in AREDS is listed in the right column. The left panel shows the corresponding estimated coefficients. Light blue represents negative values of the estimated coefficients, representing an adverse subgroup; orange represents positive values of the estimated coefficients, representing a beneficial subgroup. $\alpha$ was set as 0.01 without Bonferroni corrections for generating candidate subgroups in gradient boosting CTree.
In Figure 6, the top-ranked subgroup exhibits a substantially higher subgroup importance score than all other selected subgroups. In the AREDS trial, this top beneficial subgroup, defined as “$rs1618927=\{\text{AG},\text{GG}\}\text{ or }rs1056437=\{\text{AT},\text{TT}\}$”, is the complement of its adverse counterpart, “rs1618927=AA and
rs1056437
=AA
[…]
AA
[…]
”, where the major/minor alleles for $rs1618927$ are A/G and for $rs1056437$ are A/T. This beneficial subgroup includes 651 patients (81%) in the AREDS trial. We validated this subgroup in the AREDS2 trial, where it comprises 88 patients (80%). Among these patients, 58 (66%) had positive CATE predictions. We plotted KM curves for the top beneficial subgroup versus its complementary adverse subgroup in AREDS2 (left panel of Figure 7) and conducted a log-rank test to assess the difference between the two groups. The KM curves are clearly separated, with a statistically significant difference ( $p$-value = 0.002). Validation of the other selected subgroups can be performed similarly in the AREDS2 dataset.
Figure 7: Validation results of the top one selected subgroups on AREDS2. The top one beneficial subgroup and its complementary subgroup are “$rs1618927=\{\text{AG},\text{GG}\}\text{ or }rs1056437=\{\text{AT},\text{TT}\}$” and “$rs1618927=AA\text{ and }rs1056437=AA$”, where major/minor allele of $rs$1618927 is A/G and major/minor allele of $rs$1056437 is A/T. The left panel shows KM curves on AREDS2 by the top one subgroup identified by training a DEA-learner on AREDS with $\alpha=0.01$ without Bonferroni corrections in gradient boosting CTree. $P$-values shown on the plot were calculated from log-rank tests to compare the survival curves for patients between beneficial vs adverse subgroups. The right panel shows the distributions of predicted CATEs on AREDS2 plotted by the top one subgroup. The width of the boxes represents the sample size in the beneficial and adverse subgroups. $P$-value was calculated by performing the Wilcoxon rank sum test.
6 Conclusions and Discussions
In this paper, we propose an interpretable framework for estimating CATE in survival outcomes with integrated causal subgroup discovery. The framework incorporates causal inference methods for CATE estimation, inverse probability of censoring weighting to handle right-censored data, and interpretable machine learning algorithms for subgroup identification. It offers flexibility in estimating nuisance parameters, choosing different causal inference methods for constructing pseudo-ITEs, and applying tree-based methods for subgroup generation and selection via penalty terms. We evaluated the performance of three meta-learners for constructing pseudo-ITEs, each combined with conditional inference trees to simultaneously estimate CATE and identify subgroups. Among them, the DEA-learner demonstrated the highest subgroup identification accuracy while maintaining strong predictive performance. We applied the DEA-learner-based framework to the AREDS trial and identified genetic subgroups exhibiting treatment heterogeneity, with findings validated using the independent AREDS2 dataset.
One limitation of our proposed method is that the choice of the time of interest can influence the magnitude of estimated subgroup effects. Determining a meaningful time point requires prior knowledge or preliminary analysis. Furthermore, in implementing the conditional inference tree for unbiased splitting, it is necessary to decide whether to apply multiple testing adjustments and select a significance level ($\alpha$) for the splitting criterion. Although we provide recommendations based on simulation studies in various scenarios, the choice of $\alpha$ remains subjective. A stringent $\alpha$ level may lead to sparse subgroup identification and potentially overlook meaningful subgroups, while a more lenient $\alpha$ level may yield a larger number of subgroups, some with small effect sizes. Based on our real data analyses, we observed that the top variables—those with the highest variable importance scores—are generally consistent regardless of the chosen $\alpha$ level. We also recommend conducting additional survival analyses to further evaluate the effects of identified subgroups. Finally, our framework assumes no unmeasured confounding; thus, its performance may be limited in observational settings (e.g., electronic health records) unless a comprehensive set of covariates is available. Future work will focus on extending the proposed method to accommodate observational study designs.
Acknowledgements
The author(s) disclosed receipt of the following financial support for the research, authorship and/or publication of this article: This work was funded by [details omitted for blinded peer review].
Supporting information
Supplemental material for this article is available online. The key functions for running simulations and real data analysis are available on Github and will be made available once the manuscript is accepted.
Conflict of interest
The authors declare that they have no conflict of interest.
Data Availability Statement
Both phenotype and genotype data of AREDS and AREDS2 are available from the online repository dbGap (accession: phs000001.v3.p1 and phs001039.v1.p1, respectively).